OpenStudy (anonymous):

Solve for n: 16=8n+8n

OpenStudy (anonymous):

If you have a*n+b*n you can add them together like this: (a+b)*n.

OpenStudy (anonymous):

Where a and b can be whatever number you want!

OpenStudy (anonymous):

add the 8n+8n then divided whatever number u get to 16

OpenStudy (anonymous):

16n=16 16n/16 = 16/16 n= 1

OpenStudy (anonymous):

thanks